OpenStudy (anonymous):

Triangle MNO is similar to triangle PQR. If angle N measures 85 degrees and angle P measures 75 degrees, what is the measure of angle R? 20 75 85 160

OpenStudy (anonymous):

angle M = angle P = 75 angle N = angle Q = 85 so R = 180-P-Q = 180-75-85 = 20 degrees

OpenStudy (anonymous):

thank you ^_^

OpenStudy (anonymous):

Reciprocal angles are equal to each other in similar triangles. If ABC is similar to XYZ that means angle of A = angle of X ... B = ...Y ... C = ... Z

OpenStudy (anonymous):

thank you so much now i get it :) !!